OpenStudy (anonymous):

Help Pleaseeeeeeee!! If f(x) = 5-2x and f(3k) = f(k + 1), then f(k)=

OpenStudy (anonymous):

simple. using f(3k) = f(k + 1). plug 3k adn k+1 into the first function to get 5-2(3k)=5-2(k+1). solve for k to get 1/2. plug k back in to the original function.

OpenStudy (anonymous):

get 4

OpenStudy (anonymous):

i got -2

OpenStudy (anonymous):

show me what you did

OpenStudy (anonymous):

5-6k = 5-2k-2 5-6k = 2k+3

OpenStudy (anonymous):

2 = 8k 2/8 = k 1/4 = k

OpenStudy (anonymous):

on your second step you left out a negative on the 2k, its -2k not 2k.

OpenStudy (anonymous):

instead of 5-6k = 2k+3 its 5-6k = -2k+3

OpenStudy (anonymous):

so its 1/2 then

OpenStudy (anonymous):

correct. now plug that into your original function: f(x) = 5-2x as x